OpenStudy (tomektomeknyc):

arc sin (-1/√2) = a) 45 degrees b) - 45 degrees c) 135 degrees d) 225 degrees

ganeshie8 (ganeshie8):

wat value of sin gives u \(-1/\sqrt{2}\) ?

ganeshie8 (ganeshie8):

\(\sin 45 = 1/\sqrt{2}\) so, 45 degrees is not ur answer

ganeshie8 (ganeshie8):

\(\sin (-45) = -\sin 45 = -1/\sqrt{2}\) so, -45 degrees is ur answer

ganeshie8 (ganeshie8):

see if that makes more or less sense

OpenStudy (tomektomeknyc):

Thank you

ganeshie8 (ganeshie8):

np :)